# IncomeSCM: From tabular data set to time-series simulator and causal estimation benchmark

### v2

@@ Results.
We present the results for causal effect estimation Tasks 1 & 2 in Table 2 and for Task 3 in Table 3 and Figure 2. IPW methods and S-learner (Ridge) are excluded from the latter two since these methods do not return heterogeneous effect estimates, only estimates of the ATE. The fit to observables was good for both propensity models (AUC $\approx 0.98$ in predicting observed treatments) and regression estimators ($R^{2}\approx$ 0.80 in predicting observed outcomes). Despite this, all causal estimation tasks were challenging for the methods to solve. The best methods achieved CATE $R^{2}$ of only 0.30, 0.20 and 0.33 on Tasks 1, 2 and 3, respectively, falling short of explaining the majority of variation in these effects with the respective conditioning set. This is likely in large part to the separation between intervention and outcome by a horizon of six time steps (years), which introduces substantial uncertainty and functional complexity into the process.
No method is consistently at the top: The best CATE estimator is not the best ATE estimator and the best CATE estimators are different for Tasks 1, 2 and 3. More importantly, the best factual fit (e.g., S-learner (RF) in Task 1) does not yield the best effect estimates. Despite great discrimination of the fitted propensity scores, IPW estimators generally returned ATE estimates with large error. These results show the importance of constructing several evaluation tasks to understand which methods are preferable in which circumstances. For example, in Figure 2, we see that S-learner (XGB) captures very little of the variation in CATE with education, despite passable performance in Task 1.
@@ 6 Discussion
Our benchmarking results show that models with comparable fit to observed variables can vary greatly in the quality of their causal effects estimates, highlighting the challenge of the tasks and the importance of developing purpose-built data sets, causal estimators and model selection criteria.
@@
The authors acknowledge the students of the 2023 Nordic Probabilistic AI summer school who took part in a demo of this simulator. The work was supported in part by the Wallenberg AI, Autonomous systems and Software Program (WASP), funded by the Knut & Alice Wallenberg Foundation.
@@ Appendix C Data set details
Appendix D Simulator details
D.1 Causal graph
@@
Appendix E Estimator details

### v3

@@ Results.
We present the results for causal effect estimation Tasks 1 & 2 in Table 2 and for Task 3 in Table 3 and Figure 2. IPW methods and S-learner (Ridge) are excluded from the latter two since these methods do not return heterogeneous effect estimates, only estimates of the ATE. The fit to observables was good for both propensity models (AUC $\approx 0.98$ in predicting observed treatments) and regression estimators ($R^{2}\approx$ 0.80 in predicting observed outcomes). Despite this, all causal estimation tasks were challenging for the methods to solve. The best methods achieved CATE $R^{2}$ of only 0.30, 0.20, and 0.33 on Tasks 1, 2, and 3, respectively, falling short of explaining the majority of variation in these effects with the respective conditioning set. This is likely in large part to the separation between intervention and outcome by a horizon of six time steps (years), which introduces substantial uncertainty and functional complexity into the process.
No method is consistently at the top: The best CATE estimator is not the best ATE estimator and the best CATE estimators are different for Tasks 1, 2, and 3. More importantly, the best factual fit (e.g., S-learner (RF) in Task 1) does not yield the best effect estimates. Despite great discrimination of the fitted propensity scores, IPW estimators generally returned ATE estimates with large errors. These results show the importance of constructing several evaluation tasks to understand which methods are preferable in which circumstances. For example, in Figure 2, we see that S-learner (XGB) captures very little of the variation in CATE with education, despite passable performance in Task 1.
@@ 6 Discussion
Our benchmarking results show that models with comparable fit to observed variables can vary greatly in the quality of their causal effects estimates, highlighting the challenge of the tasks and the importance of developing purpose-built data sets, causal estimators, and model selection criteria.
@@
The authors would like to thank Anton Matsson for valuable discussions in the preparation of this manuscript and acknowledge the students of the 2023 Nordic Probabilistic AI summer school who took part in a demo of this simulator. The work was supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut & Alice Wallenberg Foundation.
Appendix
@@ Appendix C Data set details
Appendix D Commonly used benchmark data sets for causal effect estimation
See Table 5 for a summary of commonly used benchmarks for causal effect estimation.
Table 5: Commonly used benchmarks in causal effect estimation. Abbreviations: Arb. = arbitrary, Synth. = Synthetic, S-synth. = semi-synthetic, S-exp. = Semi-experimental., G.truth = Ground truth, C.graph = Causal graph, Unk. = Unknown, Rand. = Randomized, Des.=Designed, Data-d = Data-driven.
Dataset #Samples #Cov Context Outcome Treatment G.truth C.graph
IHDP 747 25 Real Synth, S-synth, Yes Unk.
Jobs 3212 17 Real Real S-exp No Unk.
Twins 71344 46 Real Real Real No Unk.
ACIC 2016 4802 58 Real Synth, Synth, Yes Unk.
ACIC 2018 100 000 178 Real Synth, Synth, Yes Rand.
IncomeSCM-v1 Arb. (50k) 11 Data-d. S-synth, Synth, Yes Des.
Appendix E Simulator details
@@
Table 7: Descriptions of handcrafted transition models. Some models (e.g., Education) have clipping operations to ensure that the variables stays in the specified range. Others have transitions that are too complex to write in this table. We refer the reader to the code for definition these.
[…]
Sex Constant: sext = sext-1
Race Constant: ract = ract-1
Native country Constant: natt = natt-1
[…]
Marital status Change marital status according to possible transitions, age and studies. E.g., “never married” to “divorced” is disallowed. The probability of marrying is reduced for unmarried people engaged in full-time studies.
Occupation High probability of staying in the same occupation except if finishing full-time studies $p=p_{\mbox{stay}}(\mathds{1}[{\tt stu}_{t-1}\neq{\tt full}]+\frac{1}{4}\mathds{1}[{\tt stu}_{t-1}={\tt full}])$; $s\sim\mathrm{Bern}(p)$; $\tilde{\tt occ}_{t}\sim\sigma(f_{\tt occ}(\mathbf{pa}({\tt occ}_{t})))$; ${\tt occ}_{t}=s\cdot{\tt occ}_{t-1}+(1-s)\tilde{\tt occ}_{t}$
Relationship Stay in relationship status with given probability $p_{\mbox{stay}}$, sample new status from initial-state model otherwise: $s\sim\mathrm{Bern}(p_{\mbox{stay}})$; $\tilde{\tt mar}_{t}\sim\sigma(f_{\tt mar}(\mathbf{pa}({\tt mar}_{t})))$; ${\tt mar}_{t}=s\cdot{\tt mar}_{t-1}+(1-s)\tilde{\tt mar}_{t}$
Hours-per-week Perturbs the weekly hours by constructing a convex combination of the previous hours with a draw from a the initial-state model of weekly hours given parents: $\tilde{\tt hou}_{t}\sim\sigma(f_{\tt hou}(\mathbf{pa}({\tt hou}_{t})))$; ${\tt hou}_{t}=\mbox{clip}(\alpha\cdot{\tt hou}_{t-1}+(1-\alpha)\tilde{\tt hou}_{t},0,7\cdot 24)$
Capital net Samples whether there are capital gains/losses at all, depending on whether there was last year. High probability of having some capital if there was capital last year. If capital, samples a noisy version of last years value with high probability, otherwise a new value from initial-state model.
Studies High probability of staying within a program (e.g., Bachelor’s) if started. Low probability of starting full-time studies if not studying previously, or if already high income ($>50$k). Zero probability of full-time studies if doctoral degree.
Income Samples a new salary from initial-state model with current-time inputs if full-time studies ended, if previously had no job, or if previous income was $<5k$. Otherwise, add a random raise to previous income depending on previous studies. Set income to 0 if current full-time studies and to 4/5ths if engaged in day course.
Appendix F Estimator details
@@
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes]
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] All instructions and scripts are in the GitHub repository.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 5
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 5
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See references to the Adult data set.
    2. (b)
      Did you mention the license of the assets? [Yes] See Appendix B.1.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] In the linked GitHub repository.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] The data are not identifiable.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]